an angle measures 12° more than 1/2 the measure of its supplement. find the measure of the angle. Needs to show work.
Answer by mananth(16949)

angle = x
supplement = 180-x
..
x= 1/2 *(180-x)+12
x=90-x/2 +12
x+x/2=102
3x/2=102
3x=204
/3
x=68 deg.
...
m.ananth@hotmail.ca